OpenStudy (anonymous):

For f(x) = 2/x^3, find and simplify the following. f(5/2)

OpenStudy (anonymous):

sorry i wrote it wrong

OpenStudy (anonymous):

\[ f(\color{blue}x)= \frac{2}{\color{blue}x^3} \implies f\left(\color{red}{\frac 5 2}\right)= \frac{2}{\left(\color{red}{\frac 5 2}\right)^3} \]

OpenStudy (anonymous):

Color coded for more medals.

OpenStudy (anonymous):

Can you simplify?

OpenStudy (anonymous):

is it 2/125/8?

OpenStudy (anonymous):

Yeah, but you have to keep going.

OpenStudy (anonymous):

i dont know how to do the next step..

OpenStudy (anonymous):

Flip the fraction and it becomes multiplication.\[ \frac{2}{\frac{125}{8}} = 2\times\frac{8}{125} \]

OpenStudy (anonymous):

is it 16/250?

OpenStudy (anonymous):

250? Where di that come from?

OpenStudy (anonymous):

2*125

OpenStudy (anonymous):

Do you remember how to multiply fractions?

OpenStudy (anonymous):

i guess not

OpenStudy (anonymous):

so it would be 16/125 im assuming?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

thanks